White Company has two departments: cutting and finishing. The company uses a job-order costing system and computes a predetermined overhead rate in each department. The cutting department bases its rate on machine-hours, and the finishing department bases its rate on direct labour cost. At the beginning of the year, the company made the following estimates:

Department
Cutting Finishing
  Direct labour-hours 6,500 34,500
  Machine-hours 48,750 4,500
  Manufacturing overhead cost $ 487,500 $ 553,500
  Direct labour cost $ 65,000 $ 307,500
Required:
1.

Compute the predetermined overhead rate to be used in each department. (Round your answers to 2 decimal places.)

      

2.

Assume that the overhead rates that you computed in part (1) are in effect. The job cost sheet for job 203, which was started and completed during the year, showed the following:

Department
Cutting Finishing
  Direct labour-hours 8 25
  Machine-hours 75 5
  Materials requisitioned $ 540 $ 350
  Direct labour cost $ 88 $ 165

Compute the total overhead cost applied to job 203. (Do not round intermediate calculations and round your answer to 2 decimal places.)

      

3.

Would you expect substantially different amounts of overhead cost to be assigned to some jobs if the company used a plantwide overhead rate based on direct labour cost, rather than using departmental rates?

Yes
No
Answer #1

1) Cutting overhead rate = 487500/48750 = 10 per machine hour

Finishing overhead rate = 553500/307500 = 180% of labor cost

2) Total manufacturing cost

Cutting Finishing Total
Direct material 540 350 890
Direct labor 88 165 253
Overhead 750 297 1047
Total 1378 812 2190

Total overhead cost applied = 750+297 = 1047

3) Yes